Please help me solve this
Three consecutive odd positive integers are such that the squares of the first two integers is greater than the square of the third by 65. Find the integers.

Answer by CubeyThePenguin(3113)

consecutive odd integers: (x-2), x, (x+2)

(x-2)^2 + x^2 = (x+2)^2 + 65
x^2 - 4x + 4 + x^2 = x^2 + 4x + 4 + 65
2x^2 - 4x + 4 = x^2 + 4x + 69
x^2 - 8x - 65 = 0
(x - 13)(x + 5) = 0
x = 13, x = -5

x is positive, so the integers are 11, 13, and 15.